OpenStudy (fools101):

State the equation of the line that passes through the point (0,-2) with a slope of 5. A) y=-2x B) y=5x-2 C) y=-2x+5 D) y=-2

OpenStudy (fools101):

please help!

OpenStudy (aroub):

y=mx+b , m=5 , (0,-2) -2=5(0)+b -2=b y=5x-2 so its B. :)

OpenStudy (fools101):

oh thank u so much!

OpenStudy (aroub):

anytime!